OpenStudy (anonymous):

Rationalize the denominator? Please explain every step in some detail. The equation is in the comments below.

OpenStudy (anonymous):

\[\frac{ 3\sqrt{2}-2\sqrt{3} }{ 3\sqrt{2}+2\sqrt{3} }\]

OpenStudy (anonymous):

sorry im taking long

OpenStudy (anonymous):

1. Multiply the numerator and denominator by the denominator radical. 2. Simplify as needed.

OpenStudy (anonymous):

hopefully that helps :)

OpenStudy (anonymous):

No it didn't really help I need a step by step walk through I already knew that is what I needed to do I just don't really know how.

OpenStudy (anonymous):

do you know what a conjugate is?

OpenStudy (anonymous):

Yes and the conjugate for the problem is \[3\sqrt{2}-2\sqrt{3}\]

OpenStudy (anonymous):

ok so you need to multiply the numerator and denominator by that. Follow the FOIL rules